### v1

# Extending 1089 attractor to any number of digits and any number of steps

ABSTRACT
The well-known “1089” trick reflects an amazing trait of digital reversal process and reminisces of a limiting attractor in dynamical systems even though it takes only two steps. It is natural to consider the situations when the number of digits is beyond three as in the original 1089 trick, as well as situations when the number of steps is beyond two. The generalization to a larger number of digits has been mostly done by Webster’s work which we will reproduce. After the two steps of the “1089” trick for any number of digits, the resulting integers, the number of which is very low, are named “Papadakis-Webster integers” (PWI). A PWI is always divisible by 99, and the resulting quotients consist of only 0’s and 1’s, which we name “Papadakis-Webster binary strings” (PWBS). Not all binary strings could be PWBS, and we define the “hairpin pairing rule” to determine if a binary string is a PWBS. To generalize 1089 trick to any number of steps, we propose a two-option iteration system named “iterative digital reversal” (IDR) suitably interweaving additions and subtractions. The simplest limiting behavior of IDR is 2-cycles. The elements in an IDR 2-cycle are all composed of repetitions of the 10(9)L89 ($L\geq 0$) motif, and are all PWIs. The lower 2-cycle elements after division of 99 belong to the subset of PWBS that are palindromic and consist of 0- and 1-blocks with a minimal length of two. IDR also has higher $p$-cycles ($p=$ 10, 12, 71) whose elements seem to contain at least one PWI. Another interesting finding about IDR is that it contains non-periodic and diverging trajectories, as the integer values grow to infinity. In these diverging trajectories, while the number of flanking digits around the middle point increases by the iteration, the middle part has an 8-cycle rhythm or signature which has been found in all diverging trajectories. Overall, the generalization of the original 1089 trick in both “space” and “time” leads to many new patterns in integers and new phenomenology in dynamics.
1 Introduction
Prof. Acheson recalled reading about the “1089 trick” (Ball, 1905) as a 10 years old from the I-SPY magazine in the 1950s’ England (Acheson, 2010): think of any integer with three digits, making sure the first and the last digits differ by at least two (e.g., 782); reverse the digits and take the absolute difference (e.g., 782-287=495); reverse that new integer and take the sum (495+594); the end result is always 1089. What causes this uniqueness in the final answer, or in a physicist’s jargon, the universality, regardless of the differences in the initial integer (Stanley, 1999; Deft, 2006)?
For those who have heard of the 1089 trick, there are new questions raised. Among them: what about an initial integer with 4-digit, 5-digit, or any number of digits – will the end result of the above two operations still be unique? What are the other integers besides 1089 that behave like 1089? What if the first digit and the last digit are equal or only differ by one? What if the two operations are extended to any number of operations?
In this paper, we try to provide a relatively complete set of answers to these questions. The extension of 1089 trick to arbitrary number of digits was essentially answered by Roger Webster, in a probably less known paper (Webster, 1995).
In an even earlier work, Constantinos Papadakis, a Greek engineer and inventor, also attempted to generalize the 1089 trick to any number of digits (Papadakis, 1982). He described the properties of what we call here Papadakis-Webster numbers and binary strings and emphasized the significance of their very low populations. Still, his proofs were somewhat amateurish and incomplete. His work is even less known because his monograph is not in English (for historical reasons, we plan to make his monograph freely available on the Internet soon, with an introduction in English).
Webster’s and Papadakis’s results will be reproduced here. In order to extend the 1089 trick to any number of steps, we first need to design a rule concerning when to use subtraction and when to use addition. Our rule is the following: if the reversed integer is smaller, subtracting the two; if the reversed integer is larger (or equal), adding the two.
An operation that is repeated an infinite number of steps can be considered as a dynamical system. For the 3-digit situation discussed above, the step after reaching 1089 is an addition according to our rule, because the reversed integer is larger, then next integer is 1089+9801=10890. The step after reaching 10890 is a subtraction because the reversed integer is smaller, then the next integer is 10890-9801= 1089, back to the original 1089. No matter how many steps we would take further, the system is settled on the 1089-10890 2-cycle “attractor”. All 3-digit integers where the first digit is larger than or equal to the last digit by 2 are part of “basin of attraction” of the 1089-10890 attractor.
Because all 3-digit integers are in the basin of attraction – none are attractor themselves (1089 and 10890 are 4 and 5-digit integers), for 1089 trick, the state space of all integers collapse to a very small subset by our operation. There is a strong constraint on what the integers in the attractor would look like. Two important constraints on the integers after two steps were discovered by Webster: (1) these are divisible by 99; (2) after divided by 99, the quotient is a binary string (consisting of 0s and 1s only). In fact, this binary string is related to the digital-borrowing sequence during the first subtraction step. We will examine to what degree these are still true when we extend the 1089 trick to any number of steps.
To our surprise, 2-cycle attractors are not the only possible end results of our operation with infinite number of steps. Other longer cycles are possible. To our surprise, there are even non-periodic (acyclic) behaviors: the integers become larger and larger as we continue to iterate the mapping, even though subtraction part of the operation is always used. There are constraints on the integers in the diverging trajectories. There are also 8-cycle rhythm or “signature” in the digital sequence of these non-attractor integers. All these results are well beyond the original 1089 trick and Webster’s work.
The paper is organized as follows: section 2 is on extending 1089 trick to larger integers with more than 3 digits, principally studying the particular properties of the (very few) end-integers of the procedure. It has two subsections. One is about the produced end-integers themselves, the Papadakis-Webster integers. The other is about the Papadakis-Webster integers (PWI) divided by 99, which always end up to integers that consist of 0’s and 1’s only, called Papadakis-Webster binary strings (PWBS). Section 3 is about extending 1089 trick to infinite number of steps which contains five subsections. The first subsection introduces the iterative digital reversal (IDR) mapping and the general description of its dynamical behavior. The second and the third subsections describe the 2-cycle attractors of IDR and its connection to the Papadakis-Webster integers. The fourth and fifth subsections are about higher cyclic attractors, and acyclic diverging trajectories. The paper ends with the Discussion section. The Appendices contain discussion on a caveat in the 1089 trick, on the requirement that the number of digits after subtraction has to remain the same; more examples of PWI not listed in the main text; and an example of a 71-cycle.
2 Extension of 1089 trick to integers with any number of digits
2.1 Introducing the Papadakis-Webster integers and Papadakis-Webster binary strings
  • Definition
    (Digital reversal): For any integers with $n+1$ digits, $D=\sum_{i=0}^{n}a_{i}10^{i}=(a_{n}a_{n-1}\cdots a_{2}a_{1}a_{0})$, its digital reverse $rev(D)$ is defined as $rev(D)\equiv\sum_{i=0}^{n}a_{n-i}10^{i}=(a_{0}a_{1}a_{2}\cdots a_{n-1}a_{n})$.
  • Definition
    (Papadakis-Webster integers (PWI) ): For a (n+1)-digit integer $D$, assume (1) $D>rev(D)$, (2) $D$ and $D-rev(D)$ have the same number of digits; then the following two steps are carried out: (1) $E=D-rev(D)$, and (2) $F=E+rev(E)$; The end result F is defined as a Papadakis-Webster integer (PWI).
Note: $a_{n}>a_{0}+1$ is a sufficient, but not necessary, condition for $D$ and $E=D-rev(D)$ having the same number of digits. When $a_{n}=a_{0}+1$, $E$ may or may not have the same number of digits as $D$, and $F$ may or may not be a PWI.
  • Definition
    (Digital borrow and carryover sequence): In subtracting $rev(D)$ from $D$, the digital borrow sequence $\{b_{i}\}$ is defined as the binary indicator: $b_{i}=1$ if the subtraction at position $i$ borrows from position $i+1$, and $b_{i}=0$ if not. Similarly, in adding $D$ and $rev(D)$, the digital carryover sequence $\{c_{i}\}$ is defined by $c_{i}$ if the summation at position $i$ is larger than 10, and $c_{i}=0$ if not. The $\{b_{i}\}$ and $\{c_{i}\}$ sequences satisfy these relations:
    $b_{i}=\left\{\begin{matrix}1&\mbox{if $a_{i}-a_{n-i}-b_{i-1}<0$}\\ 0&\mbox{if $a_{i}-a_{n-i}-b_{i-1}\geq 0$ }\\ \end{matrix}\right.$ (1)
    $c_{i}=\left\{\begin{matrix}1&\mbox{if $a_{i}+a_{n-i}+c_{i-1}\geq 10$}\\ 0&\mbox{if $a_{i}+a_{n-i}+c_{i-1}<10$ }\\ \end{matrix}\right.$ (2)
Usually $b_{i}$ and $b_{n-i}$ have different values expect for some special situations. Similarly, $c_{i}$ and $c_{n-i}$ usually should have the same value unless $a_{i}+a_{n-i}=9$.
Proposition 2.1
A Papadakis-Webster integer only contains the digit-borrowing information during the $D-rev(D)$ step, and does not contain information about the original digits $\{a_{i}\}$.
  • Proof
    It is easy to check that the following formula:
    $E$ $=$ $D-rev(D)$ (3)
    $=$ $\sum_{i=0}^{n}a_{i}10^{i}-\sum_{i=0}^{n}a_{n-i}10^{i}$
    $=$ $\sum_{i=0}^{n}(a_{i}-a_{n-i}+10{b_{i}}-b_{(i-1)})10^{i}$
    contains the terms that correspond to the necessary borrowing operations for all digits for $D-rev(D)$, as a borrowing at position $i$ will increase the value at position $i$ by 10, and at the same time, a borrowing at position $i-1$ will decrease the value at position $i$ by 1. We define $b_{(-1)}=0$. Also $b_{n}=0$ is always true because our assumption that $a_{n}>a_{0}+1$. The same assumption also ensures that the leading digit of E can not be zero. In other words, $E$ has the same length n+1 as $D$.
    Next,
    $F$ $=$ $E+rev(E)$ (4)
    $=$ $\sum_{i=0}^{n}(a_{i}-a_{n-i}+10b_{i}-b_{(i-1)})10^{i}+\sum_{i=0}^{n}(a_{n-i}-a_{i}+10b_{(n-i)}-b_{(n-i-1)})10^{i}$
    $=$ $\sum_{i=0}^{n}(10b_{i}-b_{(i-1)}+10b_{(n-i)}-b_{(n-i-1)})10^{i}$
    Since F does not contain information on $\{a_{i}\}$, but only information in digit-borrowing during $D-rev(D)$, proposition 2.1 has been proven. $\blacksquare$
Proposition 2.1 explains why information concerning the original digits $\{a_{i}\}$ has been lost, and only partial information on which digit is larger than which other is kept. This great reduction on the detailed information is the basis for universality in 1089 trick.
Theorem 2.2
A Papadakis-Webster integer (PWI) is divisible by 99, and the quotient is a binary string (digits can only be 0 and 1s).
  • Proof
    After summation, the first two terms in Eq.(4) result to a complete mutual annihilation for all indices (note the re-indexing in the summation, and $b_{(-1)}=0,b_{n}=0$) :
    $\sum_{j=1}^{n+1}10b_{(j-1)}10^{j-1}-\sum_{i=0}^{n}b_{(i-1)}10^{i}=\sum_{j=1}^{n+1}b_{(j-1)}10^{j}-\sum_{i=0}^{n}b_{(i-1)}10^{i}=0$
    Then, the next two terms in Eq.(4) can be rewritten as (again, note the re-indexing in the summation):
    $F$ $=$ $\sum_{j=-1}^{n-1}10b_{(n-j-1)}10^{j+1}-\sum_{i=0}^{n}b_{(n-i-1)}10^{i}$ (5)
    $=$ $\sum_{j=-1}^{n-1}100\cdot b_{(n-j-1)}10^{j}-\sum_{i=0}^{n}b_{(n-i-1)}10^{i}$
    $=$ $99\cdot\sum_{i=0}^{n-1}b_{(n-i-1)}10^{i}\blacksquare$
Corollary 2.3
A Papadakis-Webster integer consists of only digits 0, 1, 8, and 9s.
As can be seen from Eq.5, a Papadakis-Webster integer consists of adding 99s shifted by arbitrary number of digit positions. If two 99s are shifted by one position, their sum is 99+990=1089; if shifted by two positions 99+9900=$9$; if shifted by more than $k\geq 2$ positions, the sum is 99$(0)_{k-2}$99. In the case of adding (e.g.) three shifted 99s, (e.g.) 99+990+9900=1089 +9900= 10989, no other types of digits are created. Combining all these possibilities, a Papadakis-Webster integer only consists of 0, 1, 8 and 9s.
  • Definition
    (Papadakis-Webster binary string (PWBS)): A Papadakis-Webster binary string is the quotient of a Papadakis-Webster integer divided by 99.
Corollary 2.4
A Papadakis-Webster binary string for integer $D$ is the reverse of the digital-borrowing binary indicator sequence for $D-rev(D)$, excluding the leading digit.
This can be seen from Eq.5 that the first binary value in F/99 is $b_{0}$ for $i=n-1$, the second is $b_{1}$ for $i=n-2$, etc., and the last binary value is $b_{(n-1)}$ for $i=0$.
Corollary 2.5
For an initial integer $D$ of length $n+1$, the length of the corresponding Papadakis-Webster binary string is $n$.
This can be seen by the upper limit of summation in Eq.5 ($n-1$ instead of $n$). Specific examples can be seen at Table 1. Note that the length of the corresponding Papadakis-Webster integer is either $n+2$ or $n+1$.
Table 1 shows all Papadakis-Webster integers when the initial integer $D<10^{7}$. Besides the well known Papadakis-Webster integers 99 and 1089 when the initial integers have 2 or 3 digits, the new Papadakis-Webster integers include $9$, 10890, 10989, 99099, 109890, 109989, etc.
All PWIs when the initial integer is less than 10 millions
n+1 num PWI PWI PWBS not allowed binary strings
2 1 99 1
3 1 1089 11 10
4 3 $9$ 101 100
10890 110
10989 111
5 3 99099 1001 1000, 1010, 1011, 1100, 1101
109890 1110
109989 $1$
6 8 990099 10001 10000, 10010, 10100, 10110,
991089 10011 10111, 11000, 11001, 11101
999999 10101
1089990 11010
1090089 11011
1098900 11100
1099890 11110
1099989 11111
7 8 9900099 100001 100000, 100010, 100100, 100101,
9901089 100011 100110, 100111, 101000, 101001,
10008999 101101 101010, 101011, 101100, 101110,
10890990 110010 101111, 110000, 110001, 110100,
10891089 110011 110101, 110110, 110111, 111000,
10998900 111100 111001, 111010, 111011, 111101
10999890 111110
10999989 111111
8 21 (see appendix)
9 21 (see appendix)
10 55
11 55
12 144
13 144
$\cdots$ $\cdots$
$(n+1)$ even $F_{n+1}$
$(n+1)$ odd $F_{n}$
Table 1: All PWI and the corresponding PWBS when the starting integer’s length (n+1, for $a_{n}a_{n-1}\cdots a_{1}a_{0}$) is 1-7. Binary strings that are not PWBS are also listed (last column). The number of PWI’s as a function of $n$ follows a (partial) Fibonacci sequence (1,1,3,3,8,8,21,21…). The PWI and the corresponding PWBS for n+1=8,9 are included in the Appendix.
The number of Papadakis-Webster integers increase gradually with the number of digits of the integer. Webster shows that the number of unique Papadakis-Webster integers as a function of digit length $n+1$ is a “stepwise” Fibonacci sequence (Webster, 1995), by which we mean that the numbers are not $F_{2}=$1, $F_{3}=$2, $F_{4}=$3, $F_{5}=$5, $F_{6}=$8, $F_{7}=$13, $\cdots$, but 1,1,3,3, 8,8,21,21, $\cdots$ (see Table 1).
On one hand, there are infinite numbers of Papadakis-Webster integers, one the other hand, the percentage of Papadakis-Webster integers out of all possible integers decreases exponentially as a function of the number of digits $n$: $5^{-0.5}\phi^{n}/10^{n}\approx 6.18^{-n}/\sqrt{5}$ (where $\phi=(1+\sqrt{5})/2\approx 1.618$ is the golden ratio)
2.2 Hairpin pairing rule for Papadakis-Webster binary strings
For each Papadakis-Webster integer, we also list the corresponding Papadakis-Webster binary string in Table 1. Not all binary strings are Papadakis-Webster binary strings. For example, 10, 100, 1000, 1010, 1011, 1100, 1101, etc. In the following proposition, we establish the existence of binary strings that are not Papadakis-Webster binary strings. Note that we will still call 0’s and 1’s in a PWBS (decimal) digits, not bits which are binary digits. The reason is that PWBSs are still defined in the decimal system, not in the binary system.
Proposition 2.6
Not all binary strings are Papadakis-Webster binary string.
  • Proof
    We prove it by two counterexamples. Suppose $n$ is an even number, and the number of digits $n+1$ is odd. It means that there is a digit exactly in the center position $a_{n/2}$. We will show that $1\cdots 10\cdots$, where 1 is the digit-borrowing indicator value for $D-rev(D)$ at position $n/2-1$, and 0 is that at the position $n/2$, can not be a Papadakis-Webster binary string. Considering the following digit-borrowing pattern:
    $\begin{matrix}&a_{n}&a_{n-1}&\cdots&a_{n/2}&a_{n/2-1}&\cdots&a_{1}&a_{0}\\ -)&a_{0}&a_{1}&\cdots&a_{n/2}&a_{n/2+1}&\cdots&a_{n-1}&a_{n}\\ \hline\cr b_{i}&(0)&&\cdots&0&1&\cdots&&1\\ \end{matrix}$ (6)
    which means $a_{n/2-1}<a_{n/2+1}$ or $a_{n/2-1}-1<a_{n/2+1}$, (need to borrow) and $a_{n/2}-1\geq a_{n/2}$ (no need to borrow). But the latter inequality, implying $-1\geq 0$, is impossible. Therefore, the reverse of the digit-borrowing binary string (after removing the leading digit 0, in parenthesis), $1\cdots 10\cdots$, cannot be a Papadakis-Webster binary string.
    In the second example, suppose $n$ is odd (then the number of digits is even). The central two positions are $(n+1)/2$ and $(n-1)/2$. Considering the following digit-borrowing patterns:
    $\begin{matrix}&a_{n}&a_{n-1}&\cdots&a_{(n+1)/2}&a_{(n-1)/2}&a_{(n-3)/2}&\cdots&a_{1}&a_{0}\\ -)&a_{0}&a_{1}&\cdots&a_{(n-1)/2}&a_{(n+1)/2}&a_{(n+3)/2}&\cdots&a_{n-1}&a_{n}\\ \hline\cr b_{i}&(0)&&\cdots&0&0&1&\cdots&&1\\ \end{matrix}$ (7)
    This implies $a_{(n-1)/2}-1\geq a_{(n+1)/2}$ and $a_{(n+1)/2}\geq a_{(n-1)/2}$. Adding the two lead to $-1\geq 0$, which is impossible. Therefore, the reverse of the digit-borrowing binary sequence can not be a Papadakis-Webster binary string. $\blacksquare$
  • Definition
    (Paired positions in Papadakis-Webster binary string) Denote Papadakis-Webster binary string as $B_{n-1}B_{n-2}\cdots B_{n-k-1}\cdots B_{k-1}\cdots B_{1}B_{0}$, which is the reverse of the digit-borrowing sequence ($b_{n}b_{(n-1)}\cdots b_{1}b_{0}$), after removing $b_{n}$. The positions $k-1$ and $n-k-1(k=1,2,\cdots n-1$) are defined as paired positions.
The following graph shows the correspondence between $\{a_{i}\}$, $\{b_{i}\}$, and $\{B_{k}\}$:
$\begin{matrix}&a_{n}&a_{n-1}&\cdots&a_{n-k}&\cdots&a_{k}&\cdots&a_{1}&a_{0}\\ (-)&a_{0}&a_{1}&\cdots&a_{k}&\cdots&a_{n-k}&\cdots&a_{n-1}&a_{n}\\ \hline\cr\mbox{(digit-borrow)}&b_{n}&b_{(n-1)}&\cdots&b_{(n-k)}&\cdots&b_{k}&\cdots&b_{1}&b_{0}\\ \mbox{(reverse PWBS)}&\mbox{(removed)}&B_{0}&\cdots&B_{k-1}&\cdots&B_{n-k-1}&\cdots&B_{n-2}&B_{n-1}\\ \end{matrix}$ (8)
Note that the leading digit $B_{n-1}$ does not have a paired position.
Theorem 2.7
Papadakis-Webster binary string can not have the same value at paired positions except for two situations: the digits preceding them are both 1 and their own values are both 1, or, the digits preceding them are both 0 and their own values are both 0.
  • Proof
    Since the two paired positions are involved in the subtraction operation of $a_{k}$ and $a_{n-k}$ when the two are switched, generally speaking they can not both borrow digit, or both not borrow digit. For example, if they both borrow, and without their neighbors borrowing from them, then $a_{k}<a_{n-k}$ and $a_{n-k}<a_{k}$, which is impossible. Similarly, if they both do not borrow, whereas their neighbors borrowing from them, then $a_{k}-1\geq a_{n-k}$ and $a_{n-k}-1\geq a_{k}$, which implies $a_{k}-1\geq a_{n-k}\geq a_{k}+1$, or $-1\geq+1$, also impossible. All other combinations can be shown similarly.
    For our two exceptions, the first is equivalent to $a_{k}-1<a_{n-k}$ and $a_{n-k}-1<a_{k}$, which implies $a_{k}-1<a_{n-k}<a_{k}+1$, with a unique solution $a_{k}=a_{n-k}$. In the second exception, $a_{k}\geq a_{n-k}$ and $a_{n-k}\geq a_{k}$ also has a unique solution of $a_{k}=a_{n-k}$. $\blacksquare$
Theorem 2.7 provides an algorithm to generate all Papadakis-Webster binary strings:
  1. 1.
    Start from $B_{n-1}=1$.
  2. 2.
    Pick the next digit at $B_{n-2}$. If $B_{n-2}=1$, continue to $B_{n-3}$. If $B_{n-2}=0$, set the digit at the pairing position with $n-2$ (which is $B_{0}$) to 1.
  3. 3.
    For any k $(k=1,2,\cdots,n-1)$, when $B_{n-k-1}=0$ but $B_{n-k}=1$, set $B_{k-1}=1$; when $B_{n-k-1}=1$ but $B_{n-k}=0$, set $B_{k-1}=0$; when $B_{n-k-1}=B_{n-k}$, move to the next digit.
Note: (1) this procedure will be carried out even when the index passes the mid-point. In other words, even if the pairing position is on the left of the current position, the rule needs to be checked; (2) if the pairing position is the same as the current position, the rule still needs to be checked.
Theorem 2.7 also provides a way to check if a binary sequence is PWBS or not. Given a binary string started with 1: $x_{n-1}x_{n-2}x_{n-3}\cdots x_{1}x_{0}$; removing the leading digit $x_{n-1}=1$, then pairing the remaining digits around the middle position (pairing $x_{n-2}$ with $x_{0}$, $x_{n-3}$ with $x_{1}$, etc. The digits in the pairing position should not be the same unless that same value is 1 and the digits on their left are also 1, or, that same value is 0 and the digits on their left are also 0.
Fig.1(A)(B) show two examples. Removing the leading 1 from 10111, resulting in 0111. The middle point is the space between second 1 and the third 1. Folding the string around the middle point, with first 0 pairing with the last 1, and second 1 pairing with the third 1 (Fig.1(A). It is fine when the two pairing digits have different values, but when the second 1 is the same as the third 1, the digits on their left should also be 1s – they are not, so it is not a PWBS.
In the second example, 110010011100110110, removing the lead 1, the middle position is the 1 that separates 7 digits on the left and 7 on the right (see Fig.1(B)). The middle position pairs with itself, and their (it’s) left digit has to be 1 – indeed it is, so there is no problem. All other pairing digits have different values, and again there is no problem. Therefore, this sequence is a PWBS. Because the folding and pairing of digits in Fig.1 is reminiscent of a secondary structures of RNA (Holley et al., 1965; Grover, 2022), the hairpin or stem loop, we call Theorem 2.7 “hairpin rule for Papadakis-Webster binary strings”.
Figure 1: Illustration of the hairpin pairing rule for PWBS’s. To check if a binary string is PWBS, remove the leading 1 from the sequence, then fold the rest of the sequence around its middle position (either between two middle digits, or the middle digits itself). If all pairing digits have different values, it is a PWBS. If pairing digits are both 1’s, as long as their left digits are also both 1’s, it is fine; similarly, if pairing digits are both 0’s, it will be find if the digits on their left are also both 0’s. For the self-pairing digit in the middle position, its left digit should be the same as itself in order to be PWBS. Using this rule, (B)(C)(D) are PWBS, whereas (A) is not PWBS. The double-head arrows mark the digits in pairing position that have the same value. The single-head arrows mark the self-pairing digits.
3 Extension of 1089 trick to any number of iterations
3.1 Introducing a new iteration or mapping or dynamical system
As part of the extension of the “1089 trick”, we introduce the following iteration on integers (steps $i=1,2,\cdots\infty$) with two options:
$D_{i+1}=\left\{\begin{matrix}D_{i}-rev(D_{i})&\mbox{if $rev(D_{i})<D_{i}$}\\ D_{i}+rev(D_{i})&\mbox{if $rev(D_{i})\geq D_{i}$ }\\ \end{matrix}\right.$ (9)
where the emphasis is given to the way we transform a two-step procedure to an endless dynamical system. The original 1089 trick has a caveat that the number of digits after subtraction should not be smaller (otherwise the trick would not work, see Subsection A.1). Our iteration of Eq.9 is indifferent to this caveat, and our results are more general.
Note that the subtraction is carried out when the reverse is strictly smaller than the original integer. If the condition $<$ is changed to $\leq$, a palindrome integer will iterate to zero (and forever be zero). We keep the current conditions of Eq.9 in order to produce more interesting dynamics. To simplify the citing, we call Eq.9 “iterative digital reversal” or IDR in the remaining of the paper.
For any dynamical systems, there can be these possible dynamical behaviors:
  1. 1.
    Fixed points: $\exists I\in\mathbb{N}^{+}$ so that $\forall i>I$, $D_{i+1}=D_{i}$.
  2. 2.
    2-cycles: $\exists I\in\mathbb{N}^{+}$ so that $\forall i>I$, $D_{i+2}=D_{i}$.
  3. 3.
    Periodic with higher cycle length: $\exists I\in\mathbb{N}$ so that $\forall i>I$, $D_{i+p}=D_{i}$ ($p\in\mathbb{N}^{+}>2$).
  4. 4.
    Non-periodic: $\nexists I,p\in\mathbb{N}^{+}$, so that $D_{i+p}=D_{i}$ ($i>I$).
The item-1 in the above list, fixed point, is impossible. For IDR, if $D_{i+1}=D_{i}$, then $rev(D_{i})$=0, or $D_{i}=0$. However, we specified the conditions used in Eq.9 so that $D_{i}$ can never be zero.
Because limiting sets of a dynamical system are called “attractors”, those of our 1089-trick-inspired map, IDR, can be called “1089 attractors”, whereas the end-integers from the original 1089 trick (and its extension to any number of digits) might be considered as 1089 attractors in a narrow sense. It explains the words “1089 attractor” in our title. In the next subsection, we will study the digit patterns in the 2-cycle attractors.
3.2 Constraints on the integers in the 2-cycle attractor
Proposition 3.1
If $D_{I}$ ($I\in\mathbb{N}^{+}$) and $D_{I+1}>D_{I}$ are two integers in the limiting attractor of IDR, then $D_{I+1}=10D_{I}$.
  • Proof
    Because $D_{I+1}>D_{I}$, the second option in Eq.9 is used to map $D_{I}$ to $D_{I+1}$, and because $D_{I+2}=D_{I}<D_{I+1}$, the first option in Eq.9 is used to map $D_{I+1}$ to $D_{I+2}$:
    $D_{I+2}$ $=$ $D_{I+1}-rev(D_{I+1})$ (10)
    $=$ $D_{I}+rev(D_{I})-rev(D_{I+1})$
    Because these are 2-cycle elements, $D_{I+2}=D_{I}$, therefore $rev(D_{I})=rev(D_{{I+1}})$. There are only two possibilities: the first is that $D_{I}=D_{I+1}$, which can not be correct as there is no fixed-point solution to Eq.9 and we only consider 2-cycle here. The second possibility is that $D_{I+1}$ is $D_{I}$ following by a string of 0s. Eq.9 can only increase or decrease the length of an integer by 1, therefore, the number of trailing zeros can only be one. In other words, $D_{I+1}=D_{I}\times 10$.
Corollary 3.2
Among the two integers in the 2-cycle of limiting attractor of IDR, the subtraction and addition options in Eq.9 are alternately used.
It is because $D_{I}\rightarrow D_{I+1}$ increases the integer value so addition part of Eq.9 must be used, and $D_{I+1}\rightarrow D_{I}$ decreases the integer value, thus the subtraction is used.
Corollary 3.3
For two elements in the 2-cycle attractor of IDR, $D_{I+1}>D_{I}$, then $rev(D_{I})/D_{I}=9$.
Since $D_{I+1}=D_{I}+rev(D_{I})=D_{I}\times 10$ according to Proposition 3.1, $rev(D_{I})=9\times D_{I}$. The fact that 9801 is divisible by its digital reversal was mentioned in (Hardy, 1940).
Proposition 3.4
If $D_{I}$ ($I\in\mathbb{N}^{+}$) and $D_{I+1}>D_{I}$ are two integers in the limiting attractor of IDR, the first two digits of $D_{I}$ are 1 and 0, and the last two digits of $D_{I}$ are 8 and 9.
  • Proof
    Using the proposition 3.1 and corollary 3.2, $D_{I}+rev(D_{I})=10\times D_{I}$. If $D_{I}=\sum_{i=0}^{n}a_{i}10^{i}$, we have $\sum_{i=0}^{n}a_{i}10^{i}+\sum_{i=0}^{n}a_{n-i}10^{i}=\sum_{i=0}^{n}a_{i}10^{i+1}$. To equate the coefficients on both sides, we use the digital carryover binary sequence $\{c_{i}\}$ defined in Eq.2:
    $a_{i}+a_{n-i}-10c_{i}+c_{i-1}=a_{i-1}$ (11)
    We can write Eq.11 explicitly for the leading and trailing digits:
    $\begin{matrix}column&n+1&n&n-1&\cdots&1&0\\ \hline\cr a_{i}:&&a_{n}=1^{(1)}&a_{n-1}\in(0,1)^{(3)}=0^{(4)}&\cdots&a_{1}\in(8,7)^{(3)}=8^{(4)}&a_{0}=9^{(2)}\\ a_{n-i}:&&a_{0}=9^{(2)}&a_{1}\in(8,7)^{(3)}=8^{(4)}&\cdots&a_{n-1}\in(0,1)^{(3)}=0^{(4)}&a_{n}=1^{(1)}\\ -10c_{i}:&&-10c_{n}=-10^{(1)}&-10c_{n-1}=0^{(4)}&\cdots&-10c_{1}=0^{(3)}&-10c_{0}=-10^{(2)}\\ +c_{i-1}:&c_{n}=1^{(1)}&c_{n-1}=0^{(4)}&c_{n-2}&\cdots&c_{0}=1^{(2)}&0\\ \hline\cr a_{i-1}&a_{n}=1^{(1)}&a_{n-1}\in(0,1)^{(3)}=0^{(4)}&a_{n-2}&\cdots&a_{0}=9^{(2)}&0\\ \end{matrix}$
    We can derive $(a_{n},a_{n-1},a_{1},a_{0})=(1,0,8,9)$, as well as $(c_{n},c_{n-1},c_{1},c_{0})=(1,0,0,1)$, by the following steps (matching the superscripts above:
    • (1) (column-(n+1)) $c_{n}$ can not be zero because we know $D_{I+1}=10D_{I}$ having one more digit than $D_{I}$. Therefore $c_{n}=1$, resulting to $a_{n}=1$.
    • (2) (column-0) $a_{0}+1-10c_{0}=0$. $c_{0}$ can not be zero, because otherwise we have $a_{0}+1=0$, with a negative solution for $a_{0}$. Therefore $c_{0}=1$, resulting in $a_{0}=9$.
    • (3) (column-1 and column-n) $a_{1}+a_{n-1}=8+10c_{1}$. $c_{1}$ can not be 1, because it implies $a_{1}=a_{n-1}=9$. but from column-n, $a_{n-1}=c_{n-1}$ has only a binary value $\in(0,1)$. Therefore, $c_{1}=0$.
    • (4) (from column-(n-1) and column-1) $8-10c_{n-1}+c_{n-2}=a_{n-2}$. $c_{n-1}$ can not be 1 because it implies $8-10+c_{n-2}=a_{n-2}$, or $a_{n-2}$ being negative. Therefore, $c_{n-1}=0$; then from column-n, we have $a_{n-1}=0$. Because $a_{1}+a_{n-1}=8$, we have $a_{1}=8$. $\blacksquare$
3.3 The 10(9)L89 ($L\geq 0$) motif in 2-cycle integers and proof that the quotients dividing by 99 are Papadakis-Webster binary strings
Table 2 shows all limiting 2-cycles of IDR when the initial integers have length 1-7, as well as some examples with even larger initial integers. Not only these confirm our proposition 3.4 that the limiting 2-cycle integers start with 10 and end with 89, but there are more specific patterns. For the 2-cycle integers, there is a fundamental building block of the form 10(9)L89 where the integer $L\geq 0$: these can be a single such block, or a symmetric arrangement of multiple blocks. This can be summarized by the following proposition:
Proposition 3.5
If $M_{L}$ is of a form of 10(9)L89 (i.e., the middle 9 repeats $L\geq 0$ times), then $M_{L}$ and other symmetric forms constructed from $M_{L}$ and padding zeros: $M_{L}(0)_{K}M_{L}$ (where $K\geq 0$), or $M_{L_{1}}(0)_{H_{1}}M_{L_{2}}\cdots(0)_{K_{m}}\cdots M_{L_{2}}(0)_{K_{1}}M_{L_{1}}$ (where integers $L_{1},L_{2},\cdots K_{1},K_{2},\cdots K_{m}\geq 0$), or $M_{L_{1}}(0)_{K_{1}}M_{L_{2}}\cdots(M)_{L_{m}}\cdots M_{L_{2}}(0)_{K_{1}}M_{L_{1}}$ (where $L_{1},L_{2},\cdots L_{m},K_{1},K_{2},\cdots\geq 0$), are limiting 2-cycles of IDR.
  • Proof
    We first prove the case of $M_{L}(0)_{K}M_{L}$ when $K>0$ and $L>0$. Since the first digit is 1 and the last digit is 9, the addition is carried out first:
    $\begin{matrix}&10(9)_{L}89&(0)_{K}&10(9)_{L}89\\ +)&98(9)_{L}01&(0)_{K}&98(9)_{L}01\\ \hline\cr&109(9)_{L-1}890&(0)_{K-1}&109(9)_{L-1}890\\ \end{matrix}$ (12)
    The next step will be subtraction because the last digit is 0:
    $\begin{matrix}&109(9)_{L-1}890&(0)_{K-1}&109(9)_{L-1}890\\ -)&098(9)_{L-1}901&(0)_{K-1}&089(9)_{L-1}901\\ \hline\cr&010(9)_{L-1}989&(0)_{K-1}&010(9)_{L-1}989\\ =&10(9)_{L}89&(0)_{K}&10(9)_{L}89\end{matrix}$ (13)
When $K=0$ or/and $L=0$, it can be checked that the result remains to be correct, due to the tailing 0 and/or leading 9 from the neighboring digits of a repeating unit. Proofs for the case of other symmetric combinations of the motifs can be shown similarly.$\blacksquare$
Our numerical runs, exhaustive for up to 9-digit input integers, and then sampling several millions of randomly selected input integers of higher digital length, have convincingly indicated that only integers in the form described by Proposition 3.5 are the lower members of the limiting 2-cycle, while the second (and higher) number is invariably the tenfold multiple of the first.
Proposition 3.6
The integers described in Proposition 3.5, as well as their 10-multiples, are the only limiting 2-cycle elements of IDR.
Outline of a proof of Proposition 3.6: similar to the proof of Proposition 3.4 where we show that the property of $D_{I+1}=10D_{I}$ forces the first two digits to be 0,1, and the last digits to be 8,9, we can continue to examine the constrain towards the middle of the sequence. For example, one can show that the first three and the last three digits can either (1,0,9$\cdots$,9,8,9), or (1,0,8$\cdots$,0,8,9). The first 4 and last 4 digits in the first situation would be (1,0,9,9, $\cdots$, 9,9,8,9), and in the second situation (1,0,8,9, $\cdots$, 1,0,8,9), etc. Once the uniqueness of $M_{L}=10(9)_{L}89$ as the fundamental 2-cycle motif is established, its symmetric concatenation with padding zeros in various forms can be shown also to be 2-cycles, similar to the proof for Proposition 3.5. In the Discussion section, we will mention that alternative proofs can be found in the literature (Hoey, 1992; Webster and Williams, 2012).
Table 2 also shows that when $D_{I}$ divided by 99, the quotient is a binary string. This binary string has some particular properties: it is symmetric with respect to the center, and the length of 1-blocks or 0-blocks is at least two. We have the following proposition:
Proposition 3.7
These binary strings after multiplied by 99 are limiting 2-cycle elements of IDR: $(1)_{L}(L\geq 2)$, or $(1)_{L}(0)_{K}(1)_{L}(L,K\geq 2)$, or $(1)_{L_{1}}(0)_{K_{1}}(1)_{L_{2}}\cdots(0)_{K_{m}}\cdots(1)_{L_{2}}(0)_{K_{1}}(1)_{L_{1}}$
$(L_{1},L_{2},\cdots,K_{1},K_{2}\cdots K_{m}\geq 2)$, or $(1)_{L_{1}}(0)_{K_{1}}(1)_{L_{2}}\cdots(1)_{L_{m}}\cdots(1)_{L_{2}}(0)_{K_{1}}(1)_{L_{1}}$
$(L_{1},L_{2},\cdots L_{m},K_{1},K_{2},\cdots\geq 2)$, and these are the only form of 99-quotient of 2-cycle elements of IDR.
  • Proof
    It is not difficult to show that 10(9)L-289 = 99 ×$\times(1)_{L}$, 1089$(0)_{L-2}$1089 = 99 ×$\times 11(0)_{L}11$, and combining the two, 10(9)L-289(0)K-210(9)L89= 99 ×$\times(1)_{L}(0)_{K}(1)_{L+2}$. Since the number 9s between 10 and 89, $L-2\geq 0$, we have $L\geq 0$. Since we require the number of 0s between motifs, $K-2\geq 2$, then $K\geq 2$. Other more complicated situations can be proven similarly. $\blacksquare$
Note that the minimum 0-block length or 1-block length is 2, versus no minimum length requirement of zero-padding in Proposition 3.5.
All 2-cycle elements of IDR when the initial integer is less than 10 millions
init seq $(1)_{k}$ or
length 2-cycle attractor divided by 99 $(1)_{k1}(0)_{k}(1)_{k1}$ or
$(1)_{k1}(0)_{k2}(1)_{k}(0)_{k2}(1)_{k1}$
1 and 2 1089, 10890 11, 110 k=2
3 + 10989, 109890 111, 1110 k=3
109989, 1099890 $1$, 11110 k=4
4 + 1099989,10999890 11111, 111110 k=5
5 + 10999989, 109999890 111111, 1111110 k=6
10891089, 108910890 110011, 1100110 $k_{1}=2,k=2$
108901089,1089010890 1100011, 11000110 $k_{1}=2,k=3$
1098910989, 10989109890 11100111, 111001110 $k_{1}=3,k=2$
6 +109999989, 1099999890 1111111, 11111110 k=7
1089001089, 10890010890 11000011, 110000110 $k_{1}=2,k=4$
1099999989, 10999999890 $11$, 111111110 k=8
7 +10890001089,108900010890 110000011, 1100000110 $k_{1}=2,k=5$
10989010989,109890109890 111000111, 1110001110 $k_{1}=3,k=3$
10999999989, 109999999890 111111111, 1111111110 k=9
108900001089,1089000010890 1100000011, 11000000110 $k_{1}=2,k=6$
109999999989,1099999999890 1111111111, 11111111110 k=10
109890010989,1098900109890 1110000111, 11100001110 $k_{1}=3,k=4$
109989109989,1099891099890 1111001111, 11110011110 $k_{1}=4,k=2$
1099890109989,10998901099890 11110001111, 111100011110 $k_{1}=4,k=3$
1089000001089,10890000010890 11000000011,110000000110 $k_{1}=2,k=7$
108910891089,1089108910890 1100110011, 11001100110 $k_{1}=2,k_{2}=2,k=2$
108901098901089,1089010989010890 1100011100011,11000111000110 $k_{1}=2,k_{2}=3,k=3$
Table 2: Integers in the limiting 2-cycle of IDR when the length of the initial integers is 1-7. The “+” sign means “plus all 2-cycle attractor integers already obtained when the initial integer length is lower”. The third column shows the quotients when these 2-cycle integers are divided by 99. The last column is an attempt to summarize the pattern of the binary strings in the third column. The last two rows show examples for constructing 2-cycle integers by juxtaposition of repeated copies of previously known 2-cycle integers.
Comparing Table 1 and 2, it can be seen that not all PWI’s, in fact very few of them (those in Table 1), can be limiting 2-cycle integers (those in Table 2), due to the special requirements for 2-cycle integers (e.g. symmetric binary string). On the other hand, all 2-cycle integers (in Table 2) are PWI’s (in Table 1), even though these do not satisfied the condition (i.e., $D-rev(D)$ having the same number of digits as $D$) in the proof of PWI (proposition 2.1). For example, 10890-09801=1089 loses one digit, and according to section 2.3, 1089+9801 is not guaranteed to be PWI. We propose the following theorem:
Theorem 3.8
The binary strings described in Conjecture 3.7, i.e., symmetric arrangement of 0-block and 1-block, whose lengths are larger or equal to 2, are PWBS.
  • Proof
    We present a proof by examples. In any binary sequence of this symmetric type, because of the symmetry and independently of the specific arrangement of any considered case, removing the first leading digit 1, then folding around the middle point, one can visually see that whenever the pairing digits have the same value, their left digits also have the same value, as we may see in Fig.1(C) and (D) where two examples of such binary sequences: 1100110011, and 1111110000111111 are shown. $\blacksquare$
Theorem 3.8 shows that all limiting 2-cycle integers of IDR are PWI.
3.4 Construction of an infinite number of $p$-cycles ($p>2$)
Two-cycles are not the only limiting attractors of IDR. Other limiting cycles are relatively rare but exist. The simplest $p$-cycle for $p>2$ we have observed is a 12-cycle listed in Table 3(A). This 12-cycle can be found by starting the iteration from a very small number, $D_{0}=$158. The integers in this 12-cycle are listed in Table 3(A).
During the iteration of Eq.9, integers become 99-divisible, usually well before becoming cyclic, after the sequence of integers passes through one subtraction and one addition which led to a PWI. Once the integer becomes 99-divisible, its subsequent integers by iteration are also 99-divisible. We can argue in the following: if an integer is 9-divisible, the sum of its digits is divisible by 9. This feature is preserved by digital reversal, subtraction, and addition. Therefore, once an integer is 9-divisible, it will continue to be 9-divisible after applying Eq.9. Similarly, if an integer is 11-divisible, the sum of its digits with alternating signs, $\sum_{i=0}^{n}(-1)^{i}a_{i}$ is divisible by 11. This feature also will not be affected by digital reversal, subtraction, and addition. Combining the two, the 99-divisibility is preserved by Eq.9.
However, 99-divisibility do not necessarily imply that these integers are PWI. Among the 12-cycle elements in Table 3(A), only 4 out of 12 quotient after dividing 99 (“99-quotient”) are binary sequences, and only 3 of them are PWBS, confirmed by theorem 2.7. There are no more common factors besides 99, though eleven out of 12 integers are divisible by 11$\times$99.
(A) An example of a 12-cycle attractor
i 12-cycle integer divided by 99
1 99099 (PW) 1001= 7*11*13
2 198198 2002 = 2*7*11*13
3 1090089 (PW) 11011 =7*11*11*13
4 10890990 (PW) 110010 =3*10*19*193
5 981189 9911 =11*17*53
6 1962378 19822 =2*11*17*53
7 10695069 108031 =7*11*23*61
8 106754670 1078330 =10*11*9803
9 30297069 306031 =11*43*647
10 126376272 1276528 =11*16*7253
11 399049893 4030807 =11*366437
12 108900 (not PW) 1100 =10*10*11
13=1 99099 1001 =7*11*13
(B) Another 12-cycle constructed by padding two 99099 separated by four zeros
i 12-cycle integer divided by 99
1 99099000099099 (PW) 1001000001001
2 198198000198198 2002000002002
3 1090089001090089 (PW) 11011000011011
4 10890990010890990 (PW) 110010000110010
5 981189000981189 9911000009911
6 1962378001962378 19822000019822
7 10695069010695069 108031000108031
8 106754670106754670 1078330001078330
9 30297069030297069 306031000306031
10 126376272126376272 1276528001276528
11 399049893399049893 4030807004030807
12 108900000108900 (not PW) 1100000001100
14=1 99099000099099 (PW) 1001000001001
Table 3: (A) Integers in the simplest limiting 12-cycle of IDR. The third column is the quotient by dividing these integers by 99. Of the four binary string quotients, 3 are PWBS (see Table 1) and 1 is not. Further prime factorization of these 12 integers shows that besides 99, there are no other common factors. (B) Construction of another 12-cycle by concatenating the simplest 12-cycle elements in (A). Each line in (B) corresponds to a line in (A) in that it has two copies of the simpler element plus padding zeros in-between.
The next cycle length we are examining is $p=10$, shown in Table 4. Although it has a shorter cycle length than 12, it was found later in our numerical runs as it requires a much larger initial integer, which would not have been found if the initial $D_{0}$ is small. Of the 12 99-quotients of 10-cycle elements, two are binary strings, and both are PWBS.
Moreover, the 10-cycle is not one attractor but a whole class of them described by
109008910(9)L890991089 ($L=0,1,2,\cdots$). It is similar to the class of limiting 2-cycles described by 10(9)L89. We have verified for a lot of members of the family their limiting behavior, and the generality of the above statement may be shown the same way as that in subsection 3.3.
Concatenations of $p$-cycle elements by padding certain number of zeros in-between, are also $p$-cycle elements. For example, Table 3(B) shows that combining two 12-cycle elements 99099s (as seen in Table 3(A)) with four zeros between them, 99099 $0$ 99099, is a 12-cycle member. Comparing integers in Table 3(A) and (B), each row can find its correspondence in another table, though the number of zero-padding might be different (e.g., 399049893 in Table 3(A) vs. 399049893 399049893 in Table 3(B)).
The 10-cycle motif, $M_{L}$= 1090089 10(9)L89 0991089 ($L=0,1,2,\cdots$), can also be concatenated with padding zeros in-between, ML (0)K ML ($K=0,1,2,\cdots$), which will also be 10-cycles (result not shown). Similar to 2-cycles, this concatenation can be generalized to other forms, as long as the overall arrangement of $M_{L}$ is symmetric and there are enough number of zero spacers. We propose the following proposition:
Proposition 3.9
If $M$ is a $p$-cycle integer, we can use $M$ as a motif to construct other $p$-cycles, such as $M^{\prime}=M(0)_{K}M$, $M^{\prime}=M(0)_{K_{1}}M(0)_{K_{2}}M\cdots M\cdots M(0)_{K_{2}}M(0)_{K_{1}}M$, $M^{\prime}=M(0)_{K_{1}}M(0)_{K_{2}}M\cdots(0)_{K_{m}}\cdots M(0)_{K_{2}}M(0)_{K_{1}}M$, etc.
Outline of a proof of Proposition 3.9: similar to the proof of Proposition 3.5, since the concatenation is symmetrically arranged, each of the motif $M$ will follow its own $p$-cycle dynamics, and the padding zeros play the role of separating them. Therefore, $M^{\prime}$ is also a $p$-cycle element.
In all examples from Tables 3-4, as well as our numerical runs of Eq.9, there are always at least one Papadakis-Webster integer in the limiting cycle. We propose the following conjectures:
Conjecture 3.10
One of the integers in any $p$-cycle of IDR is a Papadakis-Webster integer.
An example of 10-cycle attractor
i 10-cycle integer divided by 99
1 1090089109890991089 (PW) 11011001110010011
2 10892080098910791990 110021011100109010
3 972378109902762189 9822001110128911
4 1953645319804635468 19733791109137732
5 10599009408940099059 107060701100405041
6 105698014389430198560 1067656711004345440
7 39806979406019302059 402090701070902041
8 134827370466517262952 1361892630974921848
9 394090086130590991383 3980707940713040317
10 10890991098910900890 (PW) $110$66449
11=1 1090089109890991089 11011001110010011
Table 4: Integers in one of the limiting 10-cycle of IDR. The third column is the quotient by dividing these integers by 99. Both binary quotients, are checked by Theorem 2.7 to be PWBS.
Examples include: 1089 ($p=2$), 1090089109890991089 ($p=10$), and 99099 ($p=12$). In Tables 3-4, we have also seen that not all $p$-cycle ($p>2$) integers are Papadakis-Webster integers, unlike the situation for $p=2$ in Table 2. Consider the repeated applications of subtraction and addition in Eq.9, we would expect that both reaching the cycle and within the cycle, one subtraction would be followed by an addition in the next step. Therefore, one would expect a PWI to emerge, and Conjecture 3.10 may seem to be natural. However, due to the caveat mentioned in Subsection A.1, it is not guaranteed that the condition in Definition 2.1 is satisfied. While Eq.9 always preserve 99-divisibility, it may not preserve PWI membership.
Table A.2 shows integers in a 71-cycle. We can use these integers to check the above hypothesis. Indeed, out of 71 integers in the limiting set, 12 of them after dividing 99 lead to binary strings. Of these 12 binary strings, 8 are checked to PWBS by the hairpin rule (Theorem 2.7). We can also show that (e.g.) 8820000289999602 (row 68 in Table A.2) can be used to form another integer 8820000289999602 00000 8820000289999602, with five padding zeros, is also part of a 71-cycle integer. Same conclusion can be reached for three copies of the motif 8820000289999602: 8820000289999602 00000 8820000289999602 00000 8820000289999602 (result not shown).
Although we have not observed other cycle length besides 2, 10, 12, and 71, we hypothesize that there are other cycle lengths:
Conjecture 3.11
Eq.9 has limiting cycles with cycle length longer than 71.
3.5 Non-periodic plus diverging trajectories with infinite number of steps
We first made a promise in the Introduction that when the 1089 trick is extended to any number of digits, or, to any number of steps by a subtraction-addition mixture of iterations, something similar to the 1089-magic would appear. Indeed, extending to any number of digits would lead to a very “privileged” set of Papadakis-Webster integers, whose membership is limited. Similarly, when extending the number of steps from 2 to infinity in IDR, the iteration often ends up to a 2-cycle, whose members are more restricted, to a subset of Papadakis-Webster integers. Eq.9 may also end up to a $p$-cycle ($p>2$): we have observed $p=10,12,71$ through extensive numerical runs. The integers in a $p$-cycle are still restricted, but less so that those in a 2-cycle. We have not fully characterized the feature for integers in the set of all $p$-cycles.
However, a new type of dynamical behavior appears from IDR. This is the item no.4 listed in Section 3.1: the non-periodic (acyclic) dynamics. However, unlike the chaotic dynamics in continuous nonlinear systems, the trajectories we have observed are not wandering irrationally in the $\mathbb{N^{+}}$ space; instead, they march to infinitely large integers in a regular fashion.
Our numerical run shows that the following integer would lead to a diverging/non-periodic trajectory: 10(9)n89(0)n with $n\geq 2$. This set of integers look deceptively similar to 10(9)n89 (with $n\geq 0$) for the limiting 2-cycle integers. However there are two major differences: the extra trailing zeros, and the longer padding of 0’s between 10 and 89. The trailing zeros, in particular, destroy the symmetry between 10 and 89, and the dynamics is no longer 2-cyclic.
It can be shown that 10(9)i89(0)i will map to ($\rightarrow$) 1(0)i+1 8 (9)i+1 $\rightarrow$ 10(9)i 89 (0)i+1 $\rightarrow$ 108(9)i+1 0 (9)i $\rightarrow$ 10 (9)i+2 89 (0)i $\rightarrow$ 1(0)i+1 998 (9)i+1 $\rightarrow$ 10(9)i+2 89(0)i+1 $\rightarrow$ 109 (0)i 9890 (9)i, and finally, maps to 10(9)i+289(0)i+2 (9). The above representation of these integers do not highlight the symmetry hidden in the sequence. For that, we propose the following conjecture.
Conjecture 3.12
For a diverging/non-periodic trajectory of IDR, the middle digit(s) follow a 8-cycle rhythm: 98, 08, 8, 9, 99, 99, 9, and 9.
Note that a cyclic pattern in the middle digits do not necessarily imply that the integers themselves are cyclic. In fact, the flanking digits (both left and right) change and increase in length during the iteration for diverging trajectories. To illustrate Conjecture 3.12, we rewrite the first 16 integers starting from 10998900: 10998900, 10008999, 109989000, 108999099, 1099998900, 1000998999, 10999989000, 10900989099, and 109999890000, 109900890099, 1099998900000, 1099899900099, 10999999890000, 10990099890099, 109999998900000 , 109990098900099.
4 Discussion
While the extension of 1089 trick to any number of digits is straightforward, the extension to any number of steps need more discussion. Our Eq.9 aims at applying both subtraction and addition of an integer $D$ and its digital reverse $rev(D)$. However, Eq.9 is not unique. For example, suppose we change the condition for applying subtraction from $rev(D_{i})<D_{i}$ to $rev(D_{i})\leq D_{i}$:
$D_{i+1}=\left\{\begin{matrix}D_{i}-rev(D_{i})&\mbox{if $rev(D_{i})\leq D_{i}$}\\ D_{i}+rev(D_{i})&\mbox{if $rev(D_{i})>D_{i}$ }\\ \end{matrix}\right.$ (14)
a palindromic $D_{i}$ will map to $D_{i+1}=0$ which would be a fixed point. Eq.14 will not generate anything that Eq.9 can not generate, whereas it would have more 0-fixed-points. Therefore, it is clear that Eq.9 has more complex behaviors than Eq.14.
Eq.9 provides a paradigm for generating multiple types of dynamics using a simple rule. The only factor that determine the limiting dynamics is the initial integer. An integer may have a particular digit borrowing (for subtracting its reverse, see Eq.1) and digit carryover (for adding its reverse, see Eq.2) pattern, and how these two binary sequences determine the limiting dynamical behavior is far from obvious. For a traditional dynamical system, the simplest situation is the fixed points, where the eigenvector/eigenvalues of the (transpose of) transition matrix could provide a complete solution (e.g., (Li et al., 2022)). In a similar way, the simplest situation for Eq.9 is 2-cycles, which we also know a great deal, aided by Proposition 3.1.
Adding or subtracting an integer’s digital reverse from the integer itself, especially if done in multiple or repetitive ways, is an operation leading to integers obeying to very specific structural restrictions. These include the digital composition, the appearance of specific digital motifs, as well as the divisibility properties of the outcome. Therefore, the resulting integers, either after only two steps or after infinite number of steps, could fall into a very small set in the $\mathbb{N^{+}}$ space. Indeed, the proportion of integers that are PWI decreases exponentially with the number of digits. The limiting 2-cycle elements of IDR are a subset of PWI, whose number is even fewer. Even though there is an infinite number of elements in a diverging trajectory, percentage-wise these still occupy a very small set since there are specific signature pattern in these integer sequences.
PWIs and cycle elements of IDR should not be confused with each other, even though the two are related. If we take all PWIs from Table 1 and A.1 as initial conditions for IDR, there are several possible outcomes. If the PWI is of the form 10(9)L89 or 10(9)L890, after two iterations, they are mapped to themselves, i.e., these are limiting 2-cycle elements of IDR. Otherwise, they can be mapped to other PWIs to form a 2-cycle, or other 12-cycles, $p$-cycles, or diverging trajectories. In other words, PWIs, being the end result after two steps in non-caveat situations, are not necessarily the end result after being iterated infinite number of steps.
Corollary 3.3 states that if $D_{I}$ is the smaller element in a limiting 2-cycle set of IDR, then $rev(D_{I})/D_{I}=9$. Integers with this property (digital reversal of an integer being divisible by that integer) is called “reverse divisible number” or “palintiple”, and has been investigated in multiple generations (Hardy, 1940; Beech, 1990; Webster and Williams, 2012; Holt, 2014; Kendrick, 2015; Holt, 2016). As we learned in Subsection 3.3, the 2-cycle elements all have a form of symmetric arrangement of the basic 10(9)L89 motifs. These 9-palintiples were actually independently discovered in https://oeis.org/A008919 , and the exactly identical form of 9-palintiple as our Propositions 3.5 and 3.6 was presented in (Webster and Williams, 2012).
Although all the results in this article concern decimal integers (i.e., with base 10), similar results can be obtained for non-decimal bases. For example, as already pointed out by Webster (Webster, 1995), the divisibility by 99 for decimal PWI will become divisibility (B+1)(B-1) (where B is the base) for non-decimal PWIs. For example, if B=2, non-decimal PWIs are always multiples of 3, if B=8, they are multiples of 63, etc. New cycle lengths have also been observed for non-decimal IDR (results not shown).
Many mathematical programs or calculators can not carry out arithmetics correctly for very larger integers. We have observed, for example, that the R software environment (https://www.r-project.org/) could produce incorrect output for simple division of a very large number. To aid such large number arithmetics, we provide a FORTRAN source code IDR1f.f90, detailed guidelines, and its executables in Windows environment at:
https://shorturl.at/PkwJ6
https://shorturl.at/0Qje5
https://shorturl.at/G1g1y
Finally, since a large number of results are presented here, we summarize the major ones in Table 5.
A summary of the results presented in this article
situation name description
two-step PWI multiples of 99 (=PWBS) (theorem 2.2)
PWBS reverse of the digit borrowing sequence (Eq.5)
PWBS hairpin pairing rule (theorem 2.7, Fig.1)
two-step with caveat no conclusion (appendix A.1)
any steps (IDR) 2-cycle basic motif 10(9)L89 ($L\geq 0$) (prop 3.4, 3.5)
2-cycle symmetric arrangements of basic motif (prop. 3.7)
2-cycle /99 symmetric arrangements of 1- and 0-blocks (length $\geq$2)(prop 3.8)
p-cycle p=10,12, 71
10-cycle a motif element: 1090089 10(9)L89 0991089 ($L\geq 1$) (sec 3.4)
p-cycle symmetric arrangements of p-cycle motifs with padding zeros (prop 3.9)
diverging 8-cycle rhythm in the middle: 98,08,8,9,99,99,9,9 (conj 3.12)
Table 5: A summary of results presented in this article.
Acknowledgement: We would like to thank Dr. Astero Provata for helpful discussions and reading the manuscript. YA thanks Dr. Marianna Vasilakaki for computer assistance, and Ms. Maria Metaxa for sharing a copy of C. Papadakis’ monograph. YA would like to dedicate this work to the memory of Michael Rossos – a friend, an old gentleman, a polymath with a special interest in linguistics and ancient religions. It was him, after reading a mathematics popularization article about Papadakis’ work, who started to experiment with numbers adding or subtracting them after reversion, conditioning on their relative values. He was already amazed by how often number 1089 appeared in the output. Through his playful creative activity, he was the true inventor of IDR almost 30 years ago.
Appendices
Appendix A.1 A caveat in applying the 1089 trick
While the condition $a_{n}>a_{0}+1$ in 1089 trick is sufficient to ensure that the integer after the subtraction step, $E$, has the same number of digits as the starting integer $D$, the condition $a_{n}=a_{0}+1$ is not sufficient. Let’s illustrate this point by the following example: if $D=4193$, where $a_{n}$ is equal to $a_{0}+1$ E=4193-3914=279 has one less digit than $D$. If we treat the leading 0 as a space-holding digit, $F=0279+9720=9999$ is indeed a Papadakis-Webster integer. However, most people would consider rev(279)=972, then $F=279+972=1351$ is no longer a Papadakis-Webster integer.
It can be seen from Eq.3 that $E$ may lose the leading digit when $a_{n}-a_{0}-b_{(n-1)}$=0 (if we exclude the situation of $a_{n}=a_{0}$, then $a_{n}=a_{0}+1$). Then the upper limit of the summation in Eq.4 is changed from $n$ to $n-1$, and we have:
$F$ $=$ $E+rev(E)$
$=$ $\sum_{i=0}^{n-1}(a_{i}-a_{n-i}+10b_{i}-b_{(i-1)})10^{i}+\sum_{i=0}^{n-1}(a_{n-i-1}-a_{i+1}+10b_{(n-i-1)}-b_{(n-i-2)})10^{i}$
Since no $a_{i}$ terms are canceled, $F$ keeps more information about the original sequence $\{a_{i}\}$. The property of uniqueness is lost, and there is no 1089 trick.
Appendix A.2 Papadakis-Webster integers from the initial integers of length 8 and 9
All PWIs when the initial integer is between 10 millions and 1 billion
n+1 PWI PWBS not allowed binary strings
8 99000099 1000001 1000000, 1000010, 1000100, 1000101,
99001089 1000011 1000110, 1001000, 1001010, 1001100,
99010989 1000111 1001101, 1001110, 1001111, 1010000,
99099099 1001001 1010001, 1010010, 1010011, 1010100,
99100089 1001011 1010110, 1010111, 1011000, 1011010,
$99$ 1010101 1011011, 1011100, 1011110, 1011111,
100089099 1011001 1100000, 1100001, 1100100, 1100101,
100098999 1011101 1100111, 1101001, 1101100, 1101101,
108900990 1100010 1101110, 1101111, 1110000, 1110001,
108901089 1100011 1110010, 1110011, 1110101, 1111001,
108910890 1100110 1111010, 1111011, 1111101
108910989 1100111
108999990 1101010
109000089 1101011
109899900 1110100
109900890 1110110
109900989 1110111
109989000 1111000
109998900 1111100
109999890 1111110
109999989 1111111
9 990000099 10000001 10000000,10000010,10000100,10000101,10000110,
990001089 10000011 10001000,10001001,10001010,10001011,10001100,
990010989 10000111 10001101,10001110,10001111,10010000,10010001,
991089099 10011001 10010010,10010011,10010100,10010101,10010110,
991090089 10011011 10010111,10011000,10011010,10011100,10011101,
999909999 10100101 10011110,10011111,10100000,10100001,10100010,
1000989099 10111001 10100011,10100100,10100110,10100111,10101000,
1000998999 10111101 10101001,$10$,10101011,10101100,10101101,
1089000990 11000010 10101110,10101111,10110000,10110001,10110010,
1089001089 11000011 10110011,10110100,10110101,10110110,10110111,
1089010890 11000110 10111000,10111010,10111011,10111100,10111110,
1089010989 11000111 10111111,11000000,11000001,11000100,11000101,
1090089990 11011010 11001000,11001001,11001010,11001011,11001100,
1090090089 11011011 11001101,11001110,11001111,11010000,11010001,
1098909900 11100100 11010010,11010011,11010100,11010101,11010110,
1098910890 11100110 11010111,11011000,11011001,11011100,11011101,
1098910989 11100111 11011110,11011111,11100000,11100001,11100010,
1099989000 11111000 11100011,11100101,11101000,11101001,11101010,
1099998900 11111100 11101011,11101100,11101101,11101110,11101111,
1099999890 11111110 11110000,11110001,11110010,11110011,11110100, 11110101,
1099999989 $11$ 11110110,11110111,11111001,11111010,11111011,11111101
Table A.1: Extension of Table 1: Papadakis-Webster integers (PWI) when the initial integers that start the two-step operation have 8 or 9 digits. The 99-quotient of the PWIs (Papadakis-Webster binary string (PWBS)) are also listed. The last column lists the binary strings that are not PWBSs.
Appendix A.3 Integers in a limiting 71-cycle
An example of a 71-cycle attractor
i 71-cycle integers divided by 99
1 9999010009899999 (not PW) 101000101110101
2 19998999010009998 202010091010202
3 109989000109999989 (PW) 1111000001111111
4 1099988901110989890 11110999001121110
5 110098790012089989 1112108990021111
6 1100079000109980000 11111909092020000
7 1099179990100279989 11102828182831111
8 10998900001099999890 (PW) 111100000011111110
9 1098900991099009989 (not PW) 11100010011101111
10 10997910893089108890 111090009021102110
11 1117712853287128989 11290028821082111
12 11015930676869306100 111272027039083900
13 10855533809265355089 109651856659246011
14 108910890100098910890 1100110001011100110
15 10891000099000891089 110010102010110011
16 108910800198000910890 1100109092909100110
17 10891799306992891089 110018174818110011
18 108911629267392610890 1100117467347400110
19 10895335504466491089 110053893984510011
20 108914801945019850890 1100149514596160110
21 10855891395911431089 109655468645570011
22 108869303355231286890 1099689932881124110
23 10187170801927318089 102900715170983011
24 108268543712734496190 1093621653663984810
25 16574106495388633389 167415217125137711
26 114907794854848780950 1160684796513624050
27 55819946396351071539 563837842387384561
28 149336961765716063394 1508454159249657206
29 642697579332885697335 6491894740736219165
30 108900991098909901089 (PW) 1100010011100100011
31 1089010900989108910890 (PW) 11000110111001100110
32 108812881099018801089 1099120011101200011
33 1088921692089207019890 10999209011002091110
34 99814662286245721089 1008228911982280011
35 1801908018019079190 18201091091101810
36 882198909910988109 8911100100110991
37 1784087929820879397 18021090200210903
38 9723868219118684268 98220891102208932
39 1099000099990000989 (PW) 11101011111010111
40 10989001099890010890 (PW) 111000011110000110
41 1187991200879911989 11999911119999111
42 11079190980901909800 111911020009110200
43 10188280071992712789 102911919919118311
44 108910009989001000890 (PW) 1100101111000010110
45 10909908999100981089 110201101001020011
46 108928809199081971990 1100291002010929010
47 9749628207173142189 98481093001748911
48 19562041924201411668 197596383072741532
49 106173452167115438259 1072459112799145841
50 1059007963928369809860 10697050140690604140
51 369918325634672800359 3736548743784573741
52 1322926602071196620322 13362894970416127478
53 3553193513773262912553 35890843573467302147
54 1000890000108999000 (not PW) 10110000001101000
55 990891990108018999 10009010001091101
56 1990702791207217098 20108109002093102
57 10897829813179288089 110079089022013011
58 108986126945072167890 1100869969142143110
59 10224856395450478089 103281377731823011
60 108312261854816320290 1094063251058750710
61 16288643396654106489 164531751481354611
62 114748789065988794750 1159078677434230250
63 57250899505000947339 578291914191928761
64 150625799555600752614 1521472722783845986
65 566882806111598278665 5726088950622204835
66 9910999989990000 (not PW) 100111111010000
67 9910000089999801 100101011010099
68 8820000289999602 89090912020198
69 6750000469999314 68181822929286
70 2610000829998738 26363644747462
71 10989000109998900 (PW) 111000001111100
72=1 9999010009899999 (not PW) 101000101110101
Table A.2: Elements in a 71-cycle of IDR. The 99-quotients of the integers on the left column is listed in the right column. If a 99-quotient is a binary sequence, we further checked if it is a PWBS or not by the hairpin pairing rule (Theorem 2.7 and Fig.1).